Two light bulbs rated at 60W/120V are connected in parallel. What is the total power used by the bulb ?
Answer #1

Total power=p1+p2(if connected in parallel)

=2*60=120 w
Answer #2


apply power = V^2/R

R = 120/(60*60)

R = 0.0333 Ohms

Power in parallel = 60*60/120

= 30 Watts